# Communication to Completion: Modeling Collaborative Workflows with Intelligent Multi-Agent Communication

### v1

@@ 4.2 Main Results
Table 2: Scalability and multi-task analysis showing communication cost and speedup across team configurations. Multi-task performance demonstrates effective workload balancing and resource sharing capabilities. Speedup is calculated as the ratio of the completion time under the "No Communication" baseline to the completion time of each respective configuration.
4.3 Scalability and Multi-Task Analysis
Table 2 demonstrates C2C’s effectiveness across team sizes and workload scenarios. The framework exhibits sub-linear scaling in communication cost: while team size increases to 3.4 times (5 to 17 agents), communication cost only increases 86%. This favorable scaling behavior stems from C2C’s intelligent communication routing and the alignment factor mechanism, which prioritizes high value interactions over communications.
Speedup analysis reveals that performance steadily improves as the number of agents increases, with the largest configuration tested (1M+16W) achieving the highest speedup of $\mathbf{1.95\times}$. This highlights the effective task parallelization and specialized skill utilization in larger teams. However, the incremental gain in speedup lessens as the team grows, suggesting that simply adding more workers yields progressively smaller advantages.
@@ 4.4 Message Type and Decomposition
Table 3 details the impact of different message types on collaboration. MEETING_INVITE messages provide the highest alignment gains (+0.27), followed by HELP_REQUEST (+0.15), directly correlating with task success by resolving critical blockers. In contrast, PROGRESS_UPDATE messages maintain synchronization with a neutral (0) impact on alignment in this context. This analysis validates C2C’s strategy of prioritizing communications that address specific knowledge gaps.
@@
Table 3: Analysis of message types in the 1M+8W configuration on Medium tasks, showing the frequency and impact of each communication type.
@@
Effective collaboration begins with high quality task decomposition. We compare C2C’s adaptive decomposition strategy against manual, naive LLM, and hierarchical methods in Table 4. C2C’s approach, which considers agent skills and workload during decomposition, achieves superior subtask clarity and higher AF. This leads to 14% increase in worker utilization rate compared to naive LLM decomposition, demonstrating that context aware planning is critical for multi-agent efficiency.
@@
Table 4: Comparison of task decomposition quality. Subtask Clarity is rated by an LLM judge (1.0 is best). C2C’s adaptive method approaches manual quality.
@@ 5 Conclusion
In this paper, we present Communication to Completion, a multi-agent framework that treats communication as an optimizable resource and quantifies task understanding through the Alignment Factor within a Sequential Action Framework. By treating alignment as a quantifiable and dynamic variable, C2C enables agents to autonomously determine when and how to communicate.
Our experiments demonstrate that this alignment-driven approach adapts naturally to task complexity, with agents communicating more frequently and through richer channels when facing difficult problems while minimizing overhead on simpler tasks. Across diverse task complexities and team sizes, C2C consistently achieves high task completion while reducing completion time and improving efficiency relative to baselines. The framework produces interpretable patterns, including alignment factor trajectories, communication type distributions, and collaboration network structures, which bridge agent behavior and coordination theory to provide design insights for multi-agent systems.
@@
While the C2C framework demonstrates significant performance gains, we acknowledge several limitations that offer avenues for future research. The findings are derived from a controlled simulation, and the framework’s performance in unpredictable, real-world environments is yet to be validated. The experiments were confined to the software engineering domain, so the C2C’s effectiveness may not generalize to other collaborative fields. Furthermore, the Alignment Factor (AF) relies on an LLM’s evaluation, which introduces potential subjectivity and bias.

### v2

@@ 4.4 Scalability and Multi-Task Analysis
Table 3: Scalability and multi-task analysis. Speedup is relative to the "No Communication" baseline.
[…]
@@ 4.5 Message Type and Decomposition
Table 4 confirms that MEETING_INVITE messages provide the highest alignment gains (+0.27), justifying their high cost for critical synchronization. HELP_REQUEST follows (+0.15), while PROGRESS_UPDATE has negligible immediate impact on AF but maintains long-term context.
@@
Table 4: Analysis of message types showing frequency and impact on AF.
@@
As shown in Table 5, C2C’s adaptive decomposition strategy achieves higher subtask clarity (0.95) compared to naive LLM prompting (0.72), approaching human manual quality (1.00).
@@
Table 5: Comparison of task decomposition quality judged by LLM.
@@ 5 Conclusion
In this paper, we investigate the modeling of collaboration efficiency and communication constraints in multi-agent systems. Rather than focusing solely on agent architecture, we propose a multi—agent framework, Communication to Completion (C2C), to quantify the interplay between information alignment and coordination costs. By formalizing task understanding through the Alignment Factor and modeling communication as a scarce resource with explicit latency, we establish a rigorous methodology for quantifying the trade-offs between coordination overhead and execution efficiency in collaborative workplace.
Our experiments validate this modeling approach, revealing that efficiency is not merely a function of individual agent capability, but a dynamic equilibrium between alignment gains and communication overhead. The simulation results demonstrate that imposing realistic constraints drives agents to emerge with strategic behaviors, such as batching questions or escalating to meetings, which are absent in unconstrained environments. These patterns hold consistent across different backbone models , suggesting that our cost-alignment model captures fundamental dynamics of collaborative intelligence.
@@
While the C2C framework demonstrates significant efficiency gains, we acknowledge several limitations that offer avenues for future research. First, the Alignment Factor acts as a proxy for understanding but does not guarantee functional correctness (e.g., passing unit tests). Second, our deterministic cost models simplify real world stochastic interruptions. Finally, our experiments are specific to software engineering; generalizing these dynamics to other collaborative domains remains future work.